OpenStudy (anonymous):

Maximize P = xy with x + y = 6.

OpenStudy (anonymous):

x + 2y = 36 x = 36 - 2y P = xy P(y) = (36-2y)y = 36y - 2y^2 P'(y) = 36 - 4y Setting this equal to 0 gives y=9 x + 2y = 36 x + 2(9) = 36 x + 18 = 36 x = 18

OpenStudy (anonymous):

Do you understand?

OpenStudy (dumbcow):

@ryanvarghese12790 , was that meant to be an example? because the constraint function doesnt match original post

OpenStudy (anonymous):

it was just an example

OpenStudy (dumbcow):

ok you may want to state that so poster is not confused

OpenStudy (anonymous):

Thank you!